What are the features of commodities suitable for using as money? What are the drawbacks...

  1.   
    1. What are the features of commodities suitable for using as money?
    2. What are the drawbacks of the following to be considered as money?

                Nails, Cigarettes, Sea Shells

  1. a. Consider the following monetary statistics.

US$

Currency held by Public                                 =             100

Reserves (Currency) held by Banks           =             300

Checking Accounts                                          =             500

Savings Accounts                                              =             1000

Estimate high powered money (monetary base), narrow money and broad money.

b. What happens to those monetary aggregates (MB, M1 and M2) in the above example, if

i. Someone withdraws $50 from a checking account, and keeps it in his wallet?

ii. Someone withdraws $50 from a checking account and deposits it back in her savings account?

  1. What is the present value of the following at the market interest rate of 10%?
    1. $500 to be received after 2 years.
    2. Income stream of $100 to be received annually during the next two years.

  1. What is the difference between a discount bond and a perpetuity?

  1. Arrange the following in ascending order based on the yields to maturity (or interest rate)
    1. Simple loan of $100 to be repaid in full after one year by paying $120.
    2. Coupon bond with face value of $1000, coupon rate of 10% remaining maturity of 6 years purchased at $900.
    3. Coupon bond with face value of $1000, coupon rate of 10% remaining maturity of 6 years purchased at $1400.
    4. Coupon bond with face value of $1000, coupon rate of 10% remaining maturity of 6 years purchased at $1000.
    1. Compare the bond prices and yields to maturity of ii, iii and iv. What are your observations?
Answer #1

A.

Commodities has different features that make it suitable for use as money. The first feature is that it does not loose its intrinsic value. The second feature is that it has the value that is free from the political impact. the third feature is that it can be stored and it derives its value, from the material from that it is made. So, commodity has real value as well. These features make commodity to be used as money.

B.

These items has following drawbacks so that they cannot be considered as money:

1. these items are perishable and not used for trading.

2. All people have no agreement upon these goods to be used as money

3. Value cannot be stored.

4. These items are not divisible in nature

-----------------

A

Monetary base = 100 + 300 = $400

Narrow money = 100 + 500 = $600

Broad money = 100+500+1000 = $1600

------------------------------------

A.

Present value = 500/1.1^2 = $413.22

B.

Present value = 100/1.1 + 100/1.1^2 = $173.55

Discount bond is sold at a discount and at maturity full face value is given, but in perpetuity bond, coupon payment takes place on annual basis till perpetuity, but face value is not paid.

Pl. repost other unanswered questions for their proper answers!